OpenStudy (anonymous):

HELP! bowling costs $2.00 a game and video games cost $.50 each. Ron bowls x games, plays y video games, and spends $22.50. Ed bowls y games, plays x video games, and spends $15.00. Write a system of equations that can be used to find how much of each game they play.

OpenStudy (anonymous):

@zepdrix sorry for bothering you soo much but can you help me again? :c

OpenStudy (australopithecus):

so first construct two equations that relate to each other based on the information you have been given

OpenStudy (australopithecus):

can you do that?

OpenStudy (anonymous):

no i need help please

OpenStudy (australopithecus):

so you cant come up with at least one equation? Hint, the second sentence gives you one of your equations

OpenStudy (anonymous):

x+y= 22.50

OpenStudy (australopithecus):

right :)

OpenStudy (anonymous):

how do i do the 2nd one?

OpenStudy (australopithecus):

now the third sentence gives you the other one

OpenStudy (anonymous):

y+x= 15.00

OpenStudy (anonymous):

but what about the other numbers? 2.00 and .50?

OpenStudy (australopithecus):

well your equations are wrong actually but they are close, now remember that x number of bowling games costs 2dollars and y number of video games cost 0.5 dollars

OpenStudy (australopithecus):

Hint you need to use multiplication

OpenStudy (anonymous):

i need help ???

OpenStudy (australopithecus):

well your equations aren't exactly wrong but, if we were to apply systems of equations to them you would get the wrong answer because you are not accounting for the cost of x or the cost of y in the equation

OpenStudy (anonymous):

ok can you tell me the equations please ?

OpenStudy (australopithecus):

if a soda costs 5 dollars and I spent 10 dollars on x number of sodas, how many sodas would I have? can you solve that?

OpenStudy (anonymous):

but i need an equation for my problem

OpenStudy (australopithecus):

I know, I'm trying to help you :) I'm not just messing with you relax this problem relates directly to your problem I pormise

OpenStudy (australopithecus):

promise*

OpenStudy (australopithecus):

Can you set up an equation for the problem I just gave you?

OpenStudy (anonymous):

i dont know how to solve it :c im not very good at it

OpenStudy (anonymous):

5x=10 ?

OpenStudy (australopithecus):

just think about it, If I spend 10 dollars on 5 dollars sodas how many would I have bought?

OpenStudy (anonymous):

2

OpenStudy (australopithecus):

right can you write an expression for that? where x = the number of sodas I purchased

OpenStudy (anonymous):

i dont know :C

OpenStudy (australopithecus):

hint use multiplication

OpenStudy (anonymous):

i really dont know :( :(

OpenStudy (australopithecus):

The expression would be 5x = 10 do you understand why?

OpenStudy (anonymous):

wait i said that above and you didnt say anything

OpenStudy (australopithecus):

lol sorry I missed that I'm making food sorry

OpenStudy (australopithecus):

Now apply that to the equations you just made, they work the same only you have two different items

OpenStudy (anonymous):

lol its ok but can you please please help me with my problem?

OpenStudy (anonymous):

i dont know how to start

OpenStudy (australopithecus):

Well if I bought two sodas one was worth 2 dollars and another was worth 3 dollars and I spent 10 dollars, could you write an expression for that?

OpenStudy (anonymous):

2x+ 3y= 10

OpenStudy (australopithecus):

:)

OpenStudy (australopithecus):

hooray

OpenStudy (australopithecus):

now do the same thing with your question

OpenStudy (anonymous):

but i dont get how to do it in my problem

OpenStudy (anonymous):

theres so many numbers

OpenStudy (australopithecus):

because it is exactly the same only with video games and bowling instead of sodas

OpenStudy (australopithecus):

well you have two different equations

OpenStudy (anonymous):

2x+0.50y= 22.50

OpenStudy (anonymous):

2x+0.50y=15 ?

OpenStudy (australopithecus):

right, now the second one

OpenStudy (anonymous):

idk theres 2 different pepple with different amounts

OpenStudy (australopithecus):

yeah perfect :)

OpenStudy (anonymous):

thats the 2 equations?

OpenStudy (australopithecus):

now for the fun part, the algebra do you understand how systems of equations work?

OpenStudy (anonymous):

yup, elimination or subsitution

OpenStudy (australopithecus):

yeah that is you are right? Do you have a better understanding why it is right now though?

OpenStudy (australopithecus):

well what you are doing really is finding where both these equations are equal

OpenStudy (anonymous):

wait how would you solve these equations

OpenStudy (australopithecus):

equal to the same thing, their intersection(s)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

but then they would both cancel out

OpenStudy (anonymous):

thats not right??

OpenStudy (australopithecus):

System of equations is easy solve for one of the variables in one of your equations

OpenStudy (australopithecus):

can you do that?

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

ok show me what you get

OpenStudy (anonymous):

y= 45+4x

OpenStudy (anonymous):

i solved for y in the 1st equation :(

OpenStudy (anonymous):

idk please help me

OpenStudy (anonymous):

????

OpenStudy (anonymous):

are you there?

OpenStudy (australopithecus):

Your two equations being: 2x+0.50y= 22.50 2x+0.50y=15 so first pick an equation and solve for one of the variables so I will pick 2x+0.50y= 22.50 now I will solve for x (it doesnt matter which variable you choose just choose the one easiest to isolate) 2x + 0.50y = 22.50 2x = 22.50 - 0.50y x = (22.50 - 0.50y)/2 x = 22.50/2 - 0.5y/2 Now it is simple as subing x into the other equation 2x+0.50y=15 so 2x+0.50y=15 becomes 2(22.50/2 - 0.5y/2) + 0.50y = 15 now you just need to solve for y once you find y you can plug that into one of your original equations and solve for x

OpenStudy (australopithecus):

yeah you are right they will cancel out ugh my brain, sorry exam time

OpenStudy (anonymous):

then the equations are wrong?? :(

OpenStudy (australopithecus):

since both lines are parallel, wow I feel dumb. https://www.wolframalpha.com/input/?i=graph+2x%2B0.50y%3D+22.50%2C+2x%2B0.50y%3D15 If you look the graphs of both of these equations you will see that they never intersect

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can you help me?

OpenStudy (australopithecus):

what systems of equations does is tells you where the two equations intersect. No the equations are correct

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

??????????

OpenStudy (anonymous):

are you there???

OpenStudy (australopithecus):

oh I just figured out the other equation 15.00 + 22.50 = 37.5 dollars total spent $37.5 = 2x + 0.5y this tells how many games in total have been played with the money spent, thus we can just use one of the two equations to figure out the amount they played each thing

OpenStudy (australopithecus):

this is a trick question, you cant even solve this problem

OpenStudy (anonymous):

yes you can theres a way idk it

OpenStudy (australopithecus):

You aren't given adequate information, like if you were told the total number of games they played you could solve it but with just this information it is unsolvable

OpenStudy (anonymous):

aw man :((((((

OpenStudy (australopithecus):

None of these equations intersect thus, you cannot apply system of equations

OpenStudy (anonymous):

wait i think u can

OpenStudy (anonymous):

they switch it up a bit: Ron bowls x games, plays y video games, and spends $22.50. Ed bowls y games, plays x video games, and spends $15.00.

OpenStudy (anonymous):

look at what x and y represent

OpenStudy (anonymous):

its different

OpenStudy (australopithecus):

you dont have adequate information sorry lol, I'm so tired and in exam study mode, I should have realized earlier. If you don't believe me you can consult someone else but I'm pretty much a 100% sure at this point.

OpenStudy (australopithecus):

you aren't given any indication of how many games they played in total thus you cannot solve this

OpenStudy (anonymous):

ok :(

OpenStudy (australopithecus):

the equations you can deduce from this have the same slope but different intercepts thus they do not intersect at any point therefore system of equations are useless in solving this. This is unsolvable

OpenStudy (anonymous):

ok thnx anyway ill ask my friend or teacher for help

OpenStudy (australopithecus):

Well they will just tell you the same thing I told you but whatevers